### v1

# Observational Signatures of Simpson-Visser Wormholes

Abstract
Observations of M87* and Sgr A* suggest that Kerr black holes can accurately model their images. However, differentiating between black holes and certain black hole mimics, such as boson stars and wormholes, remains a challenge. This study investigates the distinction between traversable Simpson-Visser wormholes and black holes through observational signatures of a luminous celestial sphere and an orbiting hot spot. We demonstrate that for a same-side light source and observer (both on the same side of the wormhole), the wormhole mimics a Schwarzschild black hole for both luminous sources. However, when the light source is positioned on the different side from the observer, photons traversing the wormhole throat generate distinct observational signatures. Specifically, the observed images are confined within a critical curve, and the light curve of the hot spot can exhibit multiple peaks with minimal centroid variations.
preprint: CTP-SCU/2024XXX
I Introduction
The groundbreaking images of the supermassive black holes candidates M87* and Sgr A*, captured by the Event Horizon Telescope (EHT) collaboration, have opened new avenues for understanding the nature in the strong field regime [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. These images reveal a characteristic feature: a dark interior region surrounded by a bright ring, which is in good agreement with theoretical predictions for Kerr black holes. This feature emerges from the strong gravitational lensing of light near light rings (or photon spheres in spherically symmetric black holes) [15, 16, 17, 18, 19, 20, 21, 22]. As a result, black hole images encode valuable information of the black hole geometry and have spurred extensive research [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43].
However, the finite resolution of the EHT observations allows alternative explanations beyond black holes. Certain horizonless ultra-compact objects (UCOs) can exhibit light rings (or photon spheres) similar to black holes, mimicking their behavior in observational simulations [44, 45, 46, 47]. Therefore, distinguishing these UCOs from black holes is a crucial topic. For example, various studies have proposed echo signals in late-time waveforms as a potential discriminant, based on the presence of a reflective surface or an extra photon sphere in specific UCO models [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. Additionally, asymmetric thin-shell wormholes with two photon spheres have been found to exhibit double shadows and an extra photon ring in their images [59, 60, 61, 62, 63]. However, the existence of multiple photon spheres outside the event horizon has also been reported for a class of hairy black holes within certain parameter spaces [64, 65, 66, 67, 68]. These multiple photon spheres in black holes can introduce features similar to those observed in wormholes, including echo signals [69], double shadows [68] and extra photon rings [70, 71]. These findings highlight the ongoing challenge of differentiating between wormholes and black holes, necessitating the development of further discriminatory methods for UCOs and black holes.
Recent observations of flaring activity near black holes, particularly the recurrent detections close to Sgr A* [72, 73, 74], have garnered significant attention. While the underlying mechanism remains unclear, it is generally believed to be attributed to magnetic reconnection within magnetized accretion disks [75, 76, 77]. Nevertheless, orbital hot spots have been employed to understand the observational signatures of these flares [78, 79, 80, 81]. Although similar investigations have been extended to hairy black holes [82] and various UCOs [83, 84], there is still a lack of research on the imaging of hot pots orbiting wormholes.
This paper investigates a static, spherically symmetric regular spacetime proposed by Simpson and Visser [85]. This spacetime interpolates between black holes, black-bounces and wormholes through a parameter denoted by $a$ Previous studies have shown that for $0<a<3M$, the regular spacetime exhibits the same shadow as a Schwarzschild black hole with identical mass and distance [86]. Notably, based on observations of the M87 galaxy’s center by the EHT, the parameter is estimated to be $a\approx 4.2M$ [87]. This implies that the wormhole can closely mimic the observational behavior of black holes. However, prior research has solely considered scenarios where the light source and observer reside on the same side of the wormhole. It is both natural and necessary to investigate scenarios where the light source and the observer are located on different sides of the wormhole. Intuitively, such a scenario would yield significantly different observational results, potentially aiding in the distinction between wormholes and black holes.
Our work focuses on the optical appearances of traversable Simpson-Visser wormholes, specifically the luminous celestial sphere and the hot spot. In Section II, we review the Simpson-Visser spacetime and discuss circular orbits for both massless and massive particles. Sections III and IV present our numerical simulations for the celestial sphere and the hot spot, respectively. Finally, we summarize our key findings in Section V. Throughout the paper, we adopt the convention $G=c=1$
II Set up
The Simpson-Visser spacetime can be described by the line element presented in [85],
$ds^{2}$ $=-f\left(r\right)dt^{2}+\frac{1}{f\left(r\right)}dr^{2}+\left(r^{2}+a^{2}\right)\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$
$f\left(r\right)$ $=1-\frac{2M}{\sqrt{r^{2}+a^{2}}},$ (1)
where $M$ is the ADM mass, and $a$ is a parameter to regularize the spacetime. The spacetime can transition from a Schwarzschild black hole to a traversable wormhole by increasing $a$ from 00. Specifically, the cases $a=0$ and $a>2M$ correspond to a Schwarzschild black hole and a traversable wormhole, respectively. Additionally, intermediate states exist, including a black-bounce for $0<a<2M$ and a one-way wormhole for $a=2M$.
The motion of test particles in the spacetime is governed by the geodesic equations
$\frac{dx^{\mu}}{d\lambda}=p^{\mu},\quad\frac{dp^{\mu}}{d\lambda}=-\Gamma_{\rho\sigma}^{\mu}p^{\rho}p^{\sigma},$ (2)
where $\lambda$ is an affine parameter, and $\Gamma_{\rho\sigma}^{\mu}$ is the Christoffel symbol. Due to the spherically symmetric and static nature of the spacetime, three conserved quantities arise that characterize the geodesics,
$E=-p_{t},\quad L_{z}=p_{\phi},\quad L^{2}=p_{\theta}^{2}+L_{z}^{2}\csc^{2}\theta.$ (3)
[…]
$\left(\frac{dr}{d\lambda}\right)^{2}+V_{\text{eff}}\left(r\right)=b^{-2},\quad V_{\text{eff}}\left(r\right)=\left(-\frac{\epsilon}{L^{2}}+\frac{1}{r^{2}+a^{2}}\right)\left(1-\frac{2M}{\sqrt{r^{2}+a^{2}}}\right),$ (4)
where $V_{\text{eff}}\left(r\right)$ is the effective potential.
Figure 1: The effective potentials for massless particles. Left: The traversable wormhole with $a=5M$ possesses a single photon sphere located at $r=0$, corresponding to a critical impact parameter $b_{\text{ph}}=5\sqrt{5/3}M$. Right: The traversable wormhole with $a=2.1M$ possesses two photon spheres located at $r_{\text{ph}}^{+}=\frac{3}{10}\sqrt{51}M$ and $r_{\text{ph}}^{-}=-\frac{3}{10}\sqrt{51}M$, corresponding to a same critical impact parameter $b_{\text{ph}}=3\sqrt{3}M$. Additionally, the wormhole possesses an anti-photon sphere at $r=0$, which however has no impact on optical observations.
In the context of massless particles, we consider a class of unstable circular orbits. These orbits correspond to a maximum point of the effective potential, denoted by
$V_{\text{eff}}\left(r_{\text{ph}}\right)=b_{\text{ph}}^{-2},\quad V_{\text{eff}}^{\prime}\left(r_{\text{ph}}\right)=0,\quad V_{\text{eff}}^{\prime\prime}\left(r_{\text{ph}}\right)<0.$ (5)
This type of orbit forms a photon sphere in the spherically symmetric spacetime, which can have a significant impact on observations. For $3M\geq a$, the wormhole possesses a single photon sphere at $r_{\text{ph}}=0$; for $2M<a<3M$, the wormhole possesses two photon spheres at $r_{\text{ph}}^{\pm}=\pm\sqrt{9M^{2}-a^{2}}$. In contrast, a minimum point of the effective potential corresponds to stable circular orbits, known as an anti-photon sphere in spherically symmetric spacetimes [88]. These anti-photon spheres have been identified in various spacetimes, including horizonless ultra-compact objects [44] and extreme static black holes [89]. Waves trapped at the anti-photon sphere can tunnel through these potential barriers and propagate either towards the event horizon or outwards to spatial infinity. This phenomenon can manifest as echo signals [69] and trigger superradiant instability [90]. However, for photons near an anti-photon sphere, potential barriers act as confining agents, rendering them invisible to distant observers and minimally affecting black hole imaging. Since the presence of a photon sphere usually greatly affects observation phenomena, we will take the wormhole of $Q=5M$ with a single photon sphere and the wormhole of $Q=2.1M$ with two photon spheres as examples to study their observational characteristics. Their effective potentials are shown in FIG. 1.
For massive particles, their stable circular orbits can usually exist within a certain region. The inner edge of this region corresponds to an innermost stable circular orbit (ISCO), satisfying the following conditions
$V_{\text{eff}}\left(r_{e}\right)=b_{e}^{-2},\quad V_{\text{eff}}^{\prime}\left(r_{e}\right)=0,\quad V_{\text{eff}}^{\prime\prime}\left(r_{e}\right)=0.$ (6)
For the traversable wormholes, the ISCOs are located at $r_{e}^{\pm}=\pm\sqrt{36M^{2}-a^{2}}$, when $2M<a<6M$. If we consider a hot spot orbiting the black hole at the ISCO on the equatorial plane, then its energy and angular momentum per unit mass can be obtained as $E_{e}=\sqrt{\frac{8}{15}}M$ and $L_{e}=\frac{6}{\sqrt{5}}M$, respectively. Consequently, the corresponding angular velocity and period are $\Omega_{e}=\frac{1}{6\sqrt{6}}M^{-1}$ and $T_{e}=12\pi\sqrt{6}M$, respectively.
To obtain the images of some light sources, we use the numerical backward ray-tracing method to calculate light rays from the observer to light sources. In the local frame of a static observer located at $\left(t_{o},r_{o},\theta_{o},\phi_{o}\right)$, the photon’s local 4-momentum $p^{(\mu)}$ can be expressed by the initial 4-momentum $p_{o}^{\mu}$,
$p^{\left(t\right)}=\frac{p_{o}^{t}}{f\left(r_{o}\right)},\quad p^{\left(r\right)}=f\left(r_{o}\right)p_{o}^{r},\quad p^{\left(\theta\right)}=\sqrt{r_{o}^{2}+a^{2}}p_{o}^{\theta},\quad p^{\left(\phi\right)}=\sqrt{r_{o}^{2}+a^{2}}\left|\sin\theta_{o}\right|p_{o}^{\phi}.$ (7)
Furthermore, considering the observation angles $\alpha$ and $\beta$ as defined in [91], the components of the local 4-momentum become
$p^{\left(r\right)}=p^{\left(t\right)}\cos\alpha\cos\beta,\quad p^{\left(\theta\right)}=p^{\left(t\right)}\sin\alpha,\quad p^{\left(\phi\right)}=p^{\left(t\right)}\cos\alpha\sin\beta.$ (8)
[…]
$x\equiv-r_{o}\beta,\quad y\equiv r_{o}\alpha.$ (9)
III Celestial Sphere
This section investigates observations of traversable wormholes illuminated by a celestial sphere. This model simulates the optical appearance of the universe as lensed by the wormhole. To illustrate the image of the celestial sphere, we position a luminous celestial sphere at $r_{\text{cs}}=50M$ or $-50M$, while a static observer is situated at $r_{o}=10M$, $\theta_{o}=\pi/2$ and $\phi_{o}=\pi$. The celestial sphere is divided into four quadrants (colored green, red, blue, and yellow) corresponding to the upper left, upper right, lower left, and lower right regions relative to the observer. Additionally, a grid of black lines is overlaid, representing lines of constant longitude and latitude, where adjacent lines are separated by $\pi/18$. For a more detailed discussion on the external view of the celestial sphere, refer to [103, 104]. To generate a simulated image, we vary the observation angles and numerically integrate the trajectories of $2000\times 2000$ photons until they intersect with the celestial sphere.
Refer to caption Refer to caption
Figure 2: The image of wormholes illuminated by the celestial sphere on the same side as the observer. The gray region represents the image of the other spacetime, bounded by the critical curve. The left panel shows the wormhole with a single photon sphere $\left(a=5M\right)$, while the right panel illustrates the case with two photon spheres $\left(a=2.1M\right)$.
We first investigate the scenario where the observer resides on the same side as the celestial sphere $\left(r_{\text{cs}}=50M\right)$, as depicted in FIG. 2. These images bear a resemblance to those of a black hole, displaying a dark region within the image confined by the critical curve. Light rays traced backward from this region would inevitably enter the other spacetime, rendering them undetectable to the observer. Notably, the size of this gray region directly correlates with the peak value of the effective potential at the photon sphere. A lower peak allows photons with larger impact parameters to pass through, resulting in a larger gray area. Furthermore, a series of compressed celestial sphere images exist outside the critical curve, asymptotically approaching it. Interestingly, wormholes with two photon spheres exhibit similar behavior to those with one, as observed photons are solely influenced by the photon sphere at $r=r_{\text{ph}}^{+}$.
Refer to caption Refer to caption
Figure 3: The image of wormholes illuminated by celestial sphere on the different side from the observer. The black region represents the image of the observer’s own spacetime, bounded by the critical curve. The left panel shows the wormhole with a single photon sphere $\left(a=5M\right)$, while the right panel illustrates the case with two photon spheres $\left(a=2.1M\right)$.
Intrigued by the possibility of observing electromagnetic radiation originating from the other side of the wormhole, we extend our analysis to scenarios where the celestial sphere resides in the other spacetime $\left(r_{\text{cs}}=-50M\right)$, as depicted in FIG. 3. The left and right panels showcase wormholes with one and two photon spheres, respectively. Light rays originating outside the critical curve cannot traverse the wormhole throat, corresponding to the observer’s spacetime (represented by the black region). Conversely, light emitted by the celestial sphere in the other spacetime can traverse the throat and be observed, forming celestial sphere images within the critical curve. Interestingly, the wormhole with two photon spheres exhibits a greater number of compressed celestial sphere images compared to the single photon sphere case. This phenomenon occurs because, in the two-photon sphere wormhole, light from the celestial sphere traverses both photon spheres at $r=r_{\text{ph}}^{-}$ and $r=r_{\text{ph}}^{+}$ before reaching the observer. As the impact parameter approaches the critical value, the increased light deflection due to the double effect of the photon spheres manifests as a more pronounced alternation of colors within the observed image.
Our analysis reveals a significant influence of the relative positioning between the celestial sphere and the observer on the resulting image. When the celestial sphere resides on the same side as the observer, the light is primarily affected by a single photon sphere, leading to an image akin to that of a Schwarzschild black hole. Conversely, when the celestial sphere is on the different side from the observer, the observed image is confined by the critical curve. In this scenario, the presence of two photon spheres significantly alters the image. Light can be concurrently influenced by both photon spheres, resulting in distinctly different images compared to the single-photon sphere case.
IV Hot Spot
This section explores observable signatures of hot spots in orbit around traversable wormholes. To simplify the analysis, we consider the hot spot as an isotropically emitting sphere with a radius of $0.25M$, following a counterclockwise path on the ISCO. Building on prior research [92, 93], observers with large incidence angles will perceive more significant Doppler effects during the hot spot motion, which can provide more information about spacetime properties. Therefore, the observers in this section are uniformly set at $\left(r_{o},\varphi_{o},\theta_{o}\right)=\left(100M,\pi,80^{\circ}\right)$. To achieve optimal precision and efficiency, we simulate the hot spot image using a $1000\times 1000$ pixel grid for each snapshot and generate $500$ snapshots for a complete orbit. Light rays are traced backward from the observer to the hot spot at each time $t_{k}$ to determine the intensity $I_{klm}$ assigned to each pixel. As outlined in [94, 83, 95, 96, 97], our analysis focuses on the following image properties:
  • Time-integrated image: This image captures the complete trajectory of the hot spot in one period by integrating the intensity over all snapshots. Mathematically, each pixel is assigned an integrated intensity,
    $\left\langle I\right\rangle_{lm}=\sum_{k}I_{klm}.$ (10)
  • Light curve: The light curve depicts the variation in the magnitude over time. The flux at each snapshot is calculated using the following formula,
    $F_{k}$ $=\sum\limits_{l}\sum\limits_{m}\Delta\Omega I_{klm}.$
    The magnitude is then derived from the flux as follows,
    $m_{k}$ $=-2.5\lg\left[F_{k}/\min\left(F_{k}\right)\right].$
  • Centroid motion: The centroid motion tracks the movement of the hot spot image. In each snapshot, the centroid is obtained by calculating the intensity-weighted average position of all pixels, normalized by the total flux,
    $\overrightarrow{c_{k}}$ $=F_{k}^{-1}\sum\limits_{l}\sum\limits_{m}\Delta\Omega I_{klm}\overrightarrow{r_{lm}},$
    Here, $\Delta\Omega$ is the solid angle per pixel, and $\overrightarrow{r_{lm}}$ is the position of the pixel relative to the image center.
These image properties, as discussed in [94, 83, 95, 96, 97], provide valuable insights into the characteristics and evolution of hot spot images within a single orbital period.
Refer to caption Refer to caption
Figure 4: The time-integrated images of hot spots orbiting wormholes on the same side as the observer, with an inclination angle of $\theta_{o}=80^{\circ}$. Pixel intensities are normalized to the maximum value for comparison. The white curve represents the critical curve, corresponding to photons with a critical impact parameter. Both the images of the single-photon sphere (Left) and double-photon sphere (Right) wormholes show two distinct image tracks.
FIG. 4 presents time-integrated images of a hot spot on the same side as the observer, captured by the observer at an inclination angle of $\theta_{o}=80^{\circ}$. Similar to observations in black holes [83, 82], both single-photon sphere and double-photon sphere wormholes exhibit two distinct image tracks. Compared to the two sub-figures, the two image tracks in the left panel are closer to the critical curve, since the wormhole with a larger $a$ has a larger critical curve.
To decipher the origin of these tracks, we introduce a numerical count, $n$, representing the number of equatorial plane crossings a light ray undergoes during its trajectory. This value uniquely characterizes light rays and the resulting image tracks. Consequently, the semicircular track corresponds to the primary image with $n=0$, while the distinct, additional track corresponds to the secondary image with $n=1$. Furthermore, the position of the hot spot relative to the wormhole dictates the location of its primary and secondary images within the tracks. When the hot spot travels in front of the wormhole, its primary image forms the lower section of the semicircle track, and its secondary image forms the upper section closer to the critical curve. Conversely, if the hot spot moves behind the wormhole, the positions of the primary and secondary images are reversed.
Both image tracks exhibit an intensity asymmetry due to the Doppler effect, where the left side is brighter than the right side. In the primary track, the brighter left side corresponds to light emitted by the hot spot when it is positioned to the left and approaching the observer. Conversely, for the secondary track, the brighter left side originates from light emitted by the hot spot when it is positioned on the right side and receding from the observer.
Figure 5: The light curve (Top) and the centroid motion (Bottom) for hot spots orbiting wormholes on the same side as the observer, with an inclination angle of $\theta_{o}=80^{\circ}$. The left and right columns correspond to the cases of $a=5M$ and $2.1M$, respectively.
Figure 6: Snapshots of same-side hot spots for the single-photon sphere (Left) and double-photon sphere (Right) wormholes, captured at the light curve peaks. The upper and lower rows depict the snapshots corresponding to the highest and second-highest peaks, respectively. The relative contribution of the $n^{\text{th}}$-order image to the total flux $F_{k}$ is given by $F_{k}^{n}/F_{k}$, where $F_{k}^{n}$ denotes the flux of the $n^{\text{th}}$-order image at $t=t_{k}$.
Our current observational capabilities are limited, allowing only for a preliminary analysis of the hot spot’s light curve and centroid motion (as depicted in FIG. 5). Notably, the light curves and centroid motions of both single-photon and double-photon sphere wormholes exhibit characteristics similar to those observed in single-photon sphere black holes. The light curve displays a prominent peak (denoted by a green dot) and a secondary, fainter peak (denoted by a blue square dot). Due to the Doppler effect, the centroid motion exhibits a leftward bias within the field of view and appears irregular due to the presence of higher-order images. These findings suggest that, based solely on hot spot observations, traversable wormholes can mimic Schwarzschild black holes. Additionally, snapshots of the light curve peak are presented in FIG. 6. These figures reveal that the primary $\left(n=0\right)$ and secondary $\left(n=1\right)$ images dominate the highest and second-highest peaks, respectively.
Refer to caption Refer to caption
Figure 7: The time-integrated images of hot spots orbiting wormholes on the different side from the observer, with an inclination angle of $\theta_{o}=80^{\circ}$. Left: The image of the single-photon sphere wormhole shows two distinct image tracks within the critical curve. Right: The image of the double-photon sphere wormhole shows four distinct image tracks within the critical curve.
We now turn to the hot spot located on the different side from the observer, orbiting at the ISCO, $r=r_{e}^{-}$. Time-integrated images for this scenario are presented in FIG. 7. Only photons with impact parameters less than the critical value, $b<b_{\text{ph}}$, can overcome the potential peak and traverse the wormhole. Consequently, the images are confined within the critical curve.
For the single-photon sphere wormhole, the hot spot image exhibits two distinct tracks. The inner and outer tracks correspond to the $n=1$ and $n=2$ images, respectively. Notably, photons traversing the wormhole throat can always cross the equatorial plane once, resulting in the absence of the $n=0$ image.
The double-photon sphere case presents a markedly different observation image compared to the single-photon case. Due to the ability of received photons to traverse two photon spheres, a stronger light deflection occurs. This phenomenon manifests as four distinct tracks in the integrated image. Ordered from the innermost outward, these tracks correspond to images with $n=1$, $2$, $3$ and $4$, respectively.
Figure 8: The light curve (Top) and the centroid motion (Bottom) for hot spots orbiting wormholes on the different side from the observer, with an inclination angle of $\theta_{o}=80^{\circ}$. The left and right columns correspond to the cases of $a=5M$ and $2.1M$, respectively.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Snapshots of different-side hot spots for wormholes with a single (Left) and double (Right) photon spheres, captured at the highest (Top), second-highest (Middle) and third-highest (Bottom) magnitude peaks.
The light curve and centroid motion of the hot spot located on the other side are presented in FIG. 8. Similar to the same-side case, the single-photon sphere wormhole exhibits a light curve with a prominent peak and a secondary peak. Consistent with this, the snapshots in the left panel of FIG. 9 reveals that the highest and second-highest peaks are dominated by the $n=1$ and $n=2$ images, respectively. Furthermore, due to image confinement within the critical curve, the centroid motion is restricted to a smaller range.
Intriguingly, the hot spot orbiting the double-photon sphere wormhole displays three distinct peaks of comparable heights in the light curve, as shown in the upper-right panel of FIG. 8. The corresponding snapshots in the right panels of FIG. 9 reveal that the highest peak is dominated by a combination of $n=2$ and $3$ images, the second-highest peak by $n=1$ and $2$ images, and the third-highest peak by $n=3$ and $4$ images. Consequently, the presence of higher-order images leads to a more irregular centroid motion in this case.
These observations of the different-side hot spot offer potential avenues for distinguishing wormholes from black holes. Notably, wormholes exhibit a more restricted range of central motion compared to black holes. Additionally, the three-peaked light curve of the double-photon sphere wormhole provides a unique signature for differentiating it from the single-photon sphere counterpart.
V Conclusions
This paper investigates the images of traversable wormholes illuminated by celestial spheres and hot spots residing in two separate universes. When the celestial sphere and hot spot reside on the same side as the observer, their images closely resemble those of a black hole. However, significant differences arise when the light sources are positioned on the different side relative to the observer:
  • Same-side scenario: The celestial sphere forms images outside the critical curve, with a sequence of compressed celestial sphere images asymptotically approaching the critical curve. The hot spot exhibits two distinct image tracks with asymmetric brightness when observed at an inclination angle of $\theta_{o}=80^{\circ}$. These two tracks lead to two peaks in the light curve.
  • Different-side scenario: The celestial sphere always forms images within the critical curve, with compressed images asymptotically approaching the critical curve. The hot spot manifests as two distinct image tracks for the single-photon sphere case, while the double-photon sphere case exhibits four distinct image tracks. In the single-photon sphere case, the two tracks again lead to two peaks in the light curve. Conversely, the combination of the four tracks in the double-photon case results in three peaks in the light curve.
By analyzing these image characteristics, we can gain valuable insights into the optical signatures of light sources near wormholes. This understanding has the potential to not only differentiate between wormholes and black holes but also discriminate between single- and double-photon sphere wormholes. While observing images from the other side of the wormhole presents an intriguing prospect, it also poses greater challenges. Therefore, higher resolution instruments, such as the next-generation Very Long Baseline Interferometry, are also keenly expected.
Acknowledgements.
We are grateful to Yiqian Chen for useful discussions and valuable comments. This work is supported in part by NSFC (Grant No. 12105191, 12275183, 12275184 and 11875196).